OpenStudy (anonymous):

what is the slope of the line (-3,-3), (1,-1)

OpenStudy (anonymous):

m=(y1-y2)/(x1-x2) so put the coordinates of first point and second point into this equation (x1=-3, y1=-3) and (x2=1, y2=-1) so slope is( -3+1)/(-3-1) =-2/-4 = 1/2

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

yes